justin, J worked 5 hours longer this week than Andy, A. They worked a total of 59 hours. Which system can be used to find out how long each worked?
A. J + 5= A
J + A=59
B. A+ 5=J
J+A= 59
C. J+ 5=A
J-A= 59
D. A+ 5= J
J-A= 59
Thank you

Answer by ikleyn(52932)

.
justin, J worked 5 hours longer this week than Andy, A. They worked a total of 59 hours.
Which system can be used to find out how long each worked?
A.  J + 5 = A
    J + A = 59

B.  A + 5 = J       <<<---+++ This system
    J + A = 59

C.  J + 5 = A
    J - A = 59

D.  A + 5 = J
    J - A = 59